# Needle: A Generative-AI Powered Monte Carlo Method for Answering Complex Natural Language Queries on Multi-modal Data

### v1

@@ 5.2.4. Anomaly Detection Threshold
6. Experiments
This section presents results of our experiments designed to evaluate the effectiveness of our proposed system, Needle. We begin by outlining the experimental setup (§ 6.1) and proceed with a proof of concept (PoC) (§ 6.2), where we demonstrate Needle ’s ability to retrieve relevant images across different scenarios.
In the first scenario, we focus on object detection tasks, which may involve identifying either rare (single occurrence) or common objects within the dataset (§ 6.2.1). In the second scenario, the goal is to locate a specific image using its natural language description. We evaluate how accurately Needle can retrieve the target tuple based on the provided description.
Following the PoC, we explore the impact of various hyper-parameters on Needle’s performance (§ 6.3) and how can Needle perform well, even with a few low-resolution guide tuples. We then analyze the performance of the anomaly detection component embedded within Needle and will compare it with human evaluation results(§ 6.4). Finally, we present a case study in which human evaluators compare Needle ’s results with CLIP (§ 6.5).
6.1. Experimental Setup
6.1.1. Datasets
We have used Caltech256 (Griffin et al., 2022), LVIS (Gupta et al., 2019), and the Berkeley DeepDrive (BDD) (Yu et al., 2020) datasets for object detection. Caltech256 is a relatively simple benchmark that includes 256 categories of objects with 30,000 images. LVIS is a long-tailed benchmark with over 1,300 categories and 100,000 images. Due to the fact that some categories in LVIS have fewer than 10 images across the entire dataset, this benchmark is considered particularly challenging for object detection tasks. BDD is a large-scale, diverse dataset for autonomous driving, which contains over 100,000 video frames and images annotated for object detection, lane detection, and drivable area segmentation.
For the targeted queries experiments, we use two image captioning benchmarks: Nocaps (Agrawal et al., 2019) and SentiCap (Mathews et al., 2016). Nocaps is a large-scale image captioning dataset that extends the Open Images dataset (Kuznetsova et al., 2020), featuring captions for novel object categories and their relations to each other. SentiCap (Mathews et al., 2016), on the other hand, focuses on sentiment-driven image captions, containing images annotated with captions that convey either positive or negative sentiments, making it suitable for evaluating models on emotional and subjective image descriptions.
6.1.2. Evaluation Metric
In our experiments, we report Average Precision (AP) as the primary evaluation metric. AP is a widely used accuracy metric in information retrieval, as it places greater emphasis on retrieving relevant results earlier in the ranking compared to simpler metrics such as precision or F1-score. AP values range from 0 to 1, with an AP of 1 indicating perfect accuracy, meaning all relevant results appear before any irrelevant ones. For a given query, AP is calculated as the average of the individual precision scores at each possible recall threshold: $AP=(\sum_{i=1}^{R}P_{i})/R$, where $R$ is the total number of relevant results in the dataset, and $P_{i}$ represents the precision when considering the top $i$ relevant results. In addition to reporting the mean Average Precision (mAP) as the final performance indicator for each baseline, we also report Hit Rate for targeted queries. Since these queries are designed to find exactly one target image, Hit Rate measures whether the relevant image appears in the results, regardless of its rank. The mean Hit Rate (mHR) across all queries is also presented for each system.
6.1.3. Baselines
In our experiments, we use the latest and highest-performing version of CLIP, specifically $ViT$-$B/32$, as the baseline model (Radford et al., 2021). CLIP (Contrastive Language–Image Pretraining) is a powerful vision-language model developed by OpenAI, designed to learn visual concepts from natural language descriptions. It achieves this by jointly training an image encoder and a text encoder using a contrastive learning objective, aligning images and their corresponding textual descriptions in a shared embedding space. CLIP has demonstrated strong zero-shot performance across a wide range of visual tasks, making it a robust baseline for image retrieval tasks.
In all experiments, we will report AP for all queries in the dataset and their “hard subset”. For each dataset, we define the hard subset as the set of tuples for which CLIP AP is less than 0.5.
6.1.4. Environmental Setup
All experiments were conducted on a local server equipped with 128 GB of memory, 32 CPU cores, and two NVIDIA GeForce RTX 2080 Ti GPUs. Four different Foundation Models were utilized throughout the experiments. runwayml-v1.5 and sdxl-turbo-v1.0 were hosted locally on the aforementioned server, while DALL·E 3 and Flux-Schnell were accessed via paid APIs.
6.2. Proof of Concepts
6.2.1. Object Detection Experiments
Table 1. Performance (mAP) Comparison Between CLIP and Needle on Object Detection Benchmark Datasets
Dataset (#hard/#all) All Queries Hard Subset
CLIP Needle CLIP Needle
BDD (2/11) 0.76 0.83 0.01 0.24
Caltech256 (10/256) 0.94 0.95 0.26 0.59
LVIS (1049/1203) 0.16 0.31 0.08 0.23
Average 0.62 0.70 0.12 0.35
((a)) LVIS

### v2

@@ C.2 Anomaly Detection in Query Images
C.3 Guide Image Caching and Metadata Generation
((a)) Caltech256
((b)) LVIS
((c)) BDD
Figure 4: Correlation between CLIP cosine similarity and Mean Average Precision (MAP) across various benchmarks. The Pearson correlation coefficient ($r$) and its associated p-value are displayed within each subplot.
The generation of guide images, while instrumental for Needle’s operation, can introduce computational overhead and latency. This is particularly true when utilizing large-scale or proprietary generative models that may incur substantial processing times and network data transmission costs. To mitigate these inefficiencies and enhance overall system performance and cost-effectiveness, Needle incorporates a novel caching mechanism designed to minimize the frequency of redundant image generation.
A fundamental challenge in open-domain text-to-image retrieval is the scarcity or inadequacy of rich, comprehensive textual metadata explicitly describing image content. As demonstrated in our experiments (Table 1 and Table 2), even advanced image tagging methods often fail to capture the full semantic depth and nuance of visual information. Ideally, the availability of precise metadata (e.g., detailed tags or captions) would enable highly efficient text-based similarity searches to retrieve relevant images directly. However, the manual generation of such high-quality metadata for large image datasets is generally infeasible.
We propose an implicit method for metadata generation and caching by leveraging the search queries themselves. The underlying premise is that for a given input query $\varphi$, any image $I_{j}$ successfully retrieved by Needle for $\varphi$ is highly likely to be semantically related to $\varphi$. This allows us to effectively ’tag’ or associate these retrieved images with the input query $\varphi$, thereby enriching our internal knowledge base. This strategy reduces the need for repeated generative processes for similar or semantically related queries, as these tagged images can then be retrieved via a similarity search solely in the text embedding space.
A critical aspect of this query-based metadata generation is quantifying the confidence in the newly established association between a retrieved image and its tagging query, especially given that the granularity of relevant objects or concepts might differ across various queries. To address this, we investigate whether the raw cosine similarity scores produced by a multi-modal embedder, such as CLIP, can serve as a reliable indicator of retrieval confidence. We hypothesize that higher similarity scores for relevant items are predictive of a more confident and accurate retrieval for a given query.
To formally validate this hypothesis, we assess the statistical correlation between an embedder’s mean cosine similarity for relevant images and its retrieval performance, as measured by Average Precision (AP). Consider a set of queries $\mathcal{Q}=\{\varphi_{k}\}_{k=1}^{N}$. For each query $\varphi_{k}$, let $\mathcal{R}_{\varphi_{k}}$ denote the set of ground-truth relevant images. The similarity between vector embeddings of an image $I$ and a query $\varphi_{k}$ using embedder $\ell$ is computed using cosine similarity, denoted as $\cos(\angle(\vec{v}_{I}^{\ell},\vec{v}_{\varphi_{k}}^{\ell}))$.
We define the mean cosine similarity (MCS) for query $q_{k}$ from an embedder $\ell$ as:
$\text{MCS}_{\ell}(\varphi_{k})=\frac{1}{|\mathcal{R}_{\varphi_{k}}|}\sum_{I\in\mathcal{R}_{\varphi_{k}}}\cos(\angle(\vec{v}_{I}^{\ell},\vec{v}_{\varphi_{k}}^{\ell}))$ (7)
We then compute the Pearson correlation coefficient, $r$, between the vector of $\{\text{MCS}_{\ell}(\varphi_{k})\}_{k=1}^{N}$ values and the vector of $\{\text{AP}(\varphi_{k})\}_{k=1}^{N}$ across all queries in $\mathcal{Q}$:
$r=\frac{\sum_{k=1}^{N}(\text{MCS}_{\ell}(\varphi_{k})-\overline{\text{MCS}})(\text{AP}(\varphi_{k})-\overline{\text{AP}})}{\sqrt{\sum_{k=1}^{N}(\text{MCS}_{\ell}(\varphi_{k})-\overline{\text{MCS}})^{2}\sum_{k=1}^{N}(\text{AP}(\varphi_{k})-\overline{\text{AP}})^{2}}}$ (8)
A positive Pearson $r$ value close to +1 indicates a strong linear relationship, suggesting that queries yielding higher mean cosine similarity with their relevant images tend to correspond to higher Average Precision. This correlation validates the utility of embedder similarity scores as a proxy for retrieval confidence. Conversely, a non-significant p-value ($p>0.05$) would imply no statistically reliable linear relationship in the observed data.
We empirically validate this correlation by calculating $r$ and its corresponding $p$-value for the Caltech256 [21], LVIS [22] and BDD100K [79] datasets using CLIP as the embedder. As illustrated in Figure 4, these experiments consistently demonstrate a strong direct correlation between the absolute value of CLIP’s cosine similarity and Mean Average Precision (MAP). This observation forms the basis of our image tagging strategy: formally, given an input query $\varphi$ and a fine-tuned threshold $\tau$, Needle finds relevant images, and in a post-retrieval stage, assigns query $\varphi$ to all images $I\in\mathcal{R}_{\varphi}$ for which $\cos(\angle(\vec{v}_{I}^{\ell},\vec{v}_{\varphi}^{\ell}))\geq\tau$. We will elaborate on how these generated tags can improve the inference speed of Needle in Appendix C.4.
C.4 Query Complexity Classifier
While Needle shows outstanding performance and accuracy in retrieval by using guide image generation in its core methodology, we observe that for many simpler queries, a large computational overhead of image generation may not be necessary to achieve satisfactory results. In such cases, computationally efficient retrieval methods, such as conventional contrastive learning approaches like CLIP [53] and ALIGN [25], might provide sufficient performance on their own.
To optimize performance and resource utilization, we propose a query complexity classifier designed to dynamically assess the complexity of an input query $\varphi$. This module is designed to predict the expected Average Precision (AP) of a query’s retrieval if performed solely using these preliminary methods. These methods, denoted as =$\mathcal{M}=\{M_{p}\}_{p=1}^{P}$, include established contrastive learning models and our proposed tag-based search mechanism (as introduced in Section C.3), all of which operate without requiring on-the-fly image generation. For an input query $\varphi$, each method $M_{p}\in\mathcal{M}$ produces a ranked list of images and their associated cosine similarity scores, denoted as $S_{M_{p}}(I,\varphi)=\cos(\angle(\vec{v}_{I},\vec{v}_{\varphi}))$. Specifically, for method $M_{p}$, the retrieval result can be represented as $R_{p}(\varphi)=\{(I_{p,j},S_{M_{p}}(I_{p,j},\varphi))\}_{j=1}^{N}$.
The design of this complexity classifier prioritizes both computational efficiency and contextual relevance. Predicting query complexity solely based on linguistic features of the query text itself can be insufficient, as a query’s actual retrieval effectiveness often depends on the specific content and nuances of the underlying image database being searched. Furthermore, for an on-the-fly decision module that must operate upon query arrival, the prediction process should be extremely fast to avoid becoming a bottleneck. To this end, we use a feature-based approach where features are derived directly from the preliminary retrieval results, which are already computed, and are designed to capture not only the query’s inherent characteristics but also its interaction with the database. This allows the classifier to operate as a lightweight model, analyzing readily available metrics rather than performing complex, potentially time-consuming, end-to-end linguistic analysis of the query text.
From the top-$K$ retrieved images and their cosine similarity scores for each preliminary method, we extract a comprehensive set of features that collectively indicate the query’s complexity. These features form a vector $\vec{f}(\varphi)\in\mathbb{R}^{D}$ and include:
  • Mean Top-$K$ Cosine Similarity Scores: For each method $M_{p}$, we compute the average cosine similarity score among its top-$K$ retrieved images. Let $R_{p,K}(\varphi)=\{I_{p,j}\}_{j=1}^{K}$ denote the set of top-$K$ images retrieved by method $M_{p}$. The mean top-$K$ cosine similarity score is:
    $\bar{S}_{M_{p},K}(\varphi)=\frac{1}{K}\sum_{I\in R_{p,K}(\varphi)}S_{M_{p}}(I,\varphi)$
    As illustrated in Figure 4, a higher $\bar{S}_{M_{p},K}(\varphi)$ generally suggests a higher expected AP for that query.
  • Top-$K$ Inter-Method Overlap Coefficients: To quantify the consensus among different preliminary methods, we compute overlap metrics between their top-$K$ retrieved sets. For any pair of distinct methods $(M_{p},M_{q})$ from $\mathcal{M}$, the Jaccard index is used:
    $J(R_{p,K}(\varphi),R_{q,K}(\varphi))=\frac{|R_{p,K}(\varphi)\cap R_{q,K}(\varphi)|}{|R_{p,K}(\varphi)\cup R_{q,K}(\varphi)|}$
    High overlap indicates strong agreement across methods, typically characteristic of queries that yield higher AP values with preliminary models.
  • Confidence Deviation: We also use measures of consistency in cosine similarity scores within the top-$K$ results for each method. The standard deviation of the top-$K$ cosine similarity scores, $\sigma_{M_{p},K}(\varphi)$, serves as an indicator of confidence deviation:
    $\sigma_{M_{p},K}(\varphi)=\sqrt{\frac{1}{K-1}\sum_{I\in R_{p,K}(\varphi)}(S_{M_{p}}(I,\varphi)-\bar{S}_{M_{p},K}(\varphi))^{2}}$
    A lower standard deviation suggests higher confidence and uniformity in cosine similarity scores for a given query, which is often observed for queries achieving higher AP.
This constructed feature vector $\vec{f}(\varphi)$ is then used as an input into a pre-trained regression model $\mathcal{C}:\mathbb{R}^{D}\to[0,1]$. For the purpose if real-time operation upon query arrival, $\mathcal{C}$ is designed as a very simple and lightweight machine learning model (e.g., a linear regression model, or a shallow neural network) that operates efficiently on the derived feature vector $\vec{f}(\varphi)$, rather than performing computationally expensive end-to-end text analysis. The regression model outputs a predicted Average Precision score, $\text{AP}_{\text{pred}}(\varphi)$, for the given query using preliminary methods. The model $\mathcal{C}$ is trained using supervised learning on a dataset of queries, where the ground-truth target for each query is its actual Average Precision (AP) score when retrieved using the preliminary methods.
A predefined threshold $\gamma_{AP}\in[0,1]$ is applied to the predicted Average Precision score $\text{AP}_{\text{pred}}(\varphi)$ to define a dynamic routing strategy:
  • If $\text{AP}_{\text{pred}}(\varphi)\geq\gamma_{AP}$, the query is categorized as "simple." In this case, Needle implements a short-circuiting mechanism, directly returning the early retrieval results obtained from the preliminary ensemble. This bypasses the computationally intensive image generation process, significantly reducing latency and computational cost.
  • If $\text{AP}_{\text{pred}}(\varphi)<\gamma_{AP}$, the query is categorized as "complex." In this scenario, Needle proceeds with its full pipeline, including guide image generation and subsequent Monte Carlo sampling, to perform a more complex retrieval.
Appendix D System Development
D.1 Development Objectives
As one of our primary objectives, we aim to develop Needle as an open-source deployment-ready application that both researchers and developers can easily adopt. Specifically, we considered a set of design objectives, outlined in the following; later, in § D.2, we describe how our system implementation fulfills these objectives.
Database Functionality and CLI Integration: Needle should be easily installable across various operating systems and offer an intuitive command-line interface (CLI) for efficient user interaction.
Deployment Readiness and Versatility: Needle should be engineered for immediate deployment, operating out-of-the-box on a wide range of system configurations, including resource-limited personal computers.
Configurability and Abstract Database Modes: Users should have the freedom to select from multiple operational modes (fast, balanced, accurate) that balance retrieval accuracy and latency according to their specific application needs. Additionally, Needle should allow for customization beyond these predefined modes.
GPU Acceleration and Scalability: Given the computational intensity of image indexing, Needle should automatically leverage available GPU resources and support workload distribution across multiple GPUs to ensure scalable performance.
Modular Embedder Architecture: As embedders are central to Needle’s functionality, the architecture should facilitate easy integration and upgrading of new embedder models, ensuring the system remains at the forefront of technological advancements.
Multi-Directory Management: Needle should be capable of managing multiple image directories concurrently. It should allow users to add, remove, or toggle directories for search while indexing new directories in the background without interrupting ongoing queries.
Robustness and Consistency Maintenance: Needle should perform consistency checks upon restart after downtime to synchronize any database changes. It should continuously monitor designated directories for modifications, ensuring the database index remains aligned with the file system.
Integration with External Image Generators: To facilitate the generation of guide images, Needle should provide connection wrappers for leading proprietary and open-source image generation services (e.g., DALL-E, Google Imagen, Replicate). Furthermore, it should allow users to integrate any image generator that complies with the specified request schema.
Flexible Output Formats and Enhanced Usability: Recognizing that Needle may serve as a component within larger pipelines, it should support multiple output formats, including JSON, YAML, and human-readable text. Additionally, built-in query previews should be provided to improve the overall user experience.
((a)) needlectl Commands Overview and Adding a Directory
((b)) Image Generator/Image Directory Configurations
((c)) Running queries using needlectl and Needle Results
((d)) Needle Query Results Builtin Preview Page for [a train] in BDD dataset [79] – 102 out of 70K images include a train
Figure 5: Screenshots of key interfaces and features of Needle.
D.2 System Details
Having defined our development objectives, we can now delve into the Needle’s implementation details.
Architecture: Needle by nature requires multiple services to work together. It needs a service for handling Image Generation, a service for handling Directory Integration and Indexing Progress, and a service for storing calculated embeddings and kNN lookup. In order to handle these multiple sources of responsibilities, we have designed Needle with a microservice architecture, delegating correlated responsibilities to their respective services. Needle’s services are:
Backend: This microservice serves as the gateway to the core functionalities of Needle. It encapsulates the primary operations of the system, exposing APIs such as /search, /directory, and /query. Additionally, it manages the initialization of embedders and oversees the image indexing process.
Database: The backend container needs persistent storage to store directory structures, indexing progress, and other critical metadata. Needle is agnostic to the choice of the database, but in the current implementation, we use PostgreSQL [^15] https://www.postgresql.org/ for this goal.
Vector Store: Needle requires a vector store that supports embedding, indexing, and inner product searches. This service is used to store the outputs of various embedders and to efficiently execute k-nearest neighbor (kNN) searches for a given query. In the current implementation, we have used Milvus [^16] https://milvus.io/ for this purpose.
Image Generator Hub: This service manages the image generation process by providing wrappers for renowned services such as DALL-E, Imagen, and Replicate. It also supports custom image generation models that conform to its defined schema. Moreover, the hub offers fallback mechanisms and priority settings to maximize robustness and ensure reliable image generation.
Command-Line Interface: Needle requires an intuitive interface to work with backend service and apply user’s requests. This service is the exposed UI of Needle to the end user.
Since the individual services have their own dependencies and requirements, we containerized all core services using docker to mitigate dependency conflicts and make deployment easier across various operating systems. We employ docker compose to orchestrate these containers, establish a shared network, and configure the individual services. Additionally, to eliminate the need for direct Docker manipulation, we developed a Command-Line Interface (CLI), needlectl, as a standalone binary. needlectl installs into the user’s path and provides seamless access to the Needle core.
Technical Stack and Frameworks: Backend and ImageGeneratorHub services built using FastAPI framework [^17] https://fastapi.tiangolo.com/. For GPU-intensive tasks, PyTorch [^18] https://pytorch.org/ is leveraged to distribute workloads across available GPUs, ensuring efficient processing.
Embedder management is streamlined via integrations with Hugging Face and the timm library [^19] https://github.com/huggingface/pytorch-image-models, facilitating the initiation and upgrading of embedding models. The CLI client (needlectl) is developed using the Typer [^20] https://typer.tiangolo.com/ library to build intuitive command-line applications, while textual [^21] https://textual.textualize.io/ is utilized to manage terminal user interfaces (TUIs).
Standalone Python libraries are packaged with PyInstaller [^22] https://pyinstaller.org/en/stable/, simplifying the distribution and execution. Installation, upgrading, and uninstallation processes are automated using Bash scripts, ensuring a smooth user experience. Containerization and orchestration are achieved with Docker [^23] https://www.docker.com/ and Docker Compose, respectively.
For documentation, we employ mdBook [^24] https://rust-lang.github.io/mdBook/ to create comprehensive and well-structured guides. Finally, all of the build, development, and versioning processes are automated via CI/CD pipelines using GitHub Actions [^25] https://github.com/features/actions.
Command-Line Interface and APIs: Needle offers a robust CLI that enables seamless interaction with its core services. The needlectl is provided as a standalone binary, allowing users to execute commands directly without requiring additional dependency installations. This design ensures that non-expert users can easily operate Needle. The CLI commands and options have been carefully designed to be intuitive and straightforward. The command structure adheres to the following format:
needlectl [--global options] [component] [action] [--action options]
This structure supports clear and consistent command usage with the following key components (some illustrative examples are provided in Listing LABEL:listing:1):
  • service: for managing Needle-related tasks
  • directory: for handling image directory operations
  • query: for managing active queries
  • generator: for interfacing with image-generation functionalities
needlectl service start     # Starts the Needle services
needlectl service log       # Retrieves service logs
needlectl directory add /path/to/image/dir --show-progress
# Adds an image directory for indexing and monitoring, displaying progress
needlectl --output json query run "a wolf" --num-engines 2 --num-images-to-generate 4 --image-quality LOW
# Executes a query using two generators, each generating four low-quality images, with JSON output
Installation and Deployment:
Needle is installed using a single one-liner command that executes a Bash script. This script automatically verifies system prerequisites and checks for GPU availability. It detects the current operating system and downloads the appropriate configuration. During installation, users are prompted to select a preferred database mode—Fast, Balanced, or Accurate—which optimizes various parameters such as the number of embedders, default image resolution for generation, default number of generators, number of images generated per query, and HNSW[42] index construction parameters (e.g., M and ef). Finally, the installation script guides users through the initial steps to start and utilize the Needle service.
Table 3: Embedder Models and Their Weights
Architecture Model Name Weight
eva[18] eva02_large_patch14_448 0.8497
regnet[75] regnety_1280 0.8235
dinoV2[51] vit_large_patch14_reg4_dinov2 0.8235
CLIP[53] vit_large_patch14_clip_336 0.8146
convnextV2[73] convnextv2_large 0.8184
bevit[7] beitv2_large_patch16_224 0.7660
Appendix E Additional Experiments
E.1 Details of Baseline Models
1. CLIP [53]: Developed by OpenAI [^26] https://huggingface.co/openai/clip-vit-base-patch32, this model uses a ViT-B/32 image encoder and a Transformer-based text encoder. It learns a shared embedding space via contrastive learning on large-scale image–text pairs, and its widespread use and benchmarking make it a standard reference. We use CLIP as our standard baseline and designate categories where CLIP’s average precision (AP) is below 0.5 as "hard categories"—instances where CLIP struggles to produce satisfactory results. This filtered subset is then used to demonstrate the enhanced capability of alternative baselines in effectively handling rare or difficult queries.
2. ALIGN [25]: Originally developed by Google, ALIGN (A Large-scale ImaGe and Noisy-text Embedding) employs a dual-encoder architecture similar to CLIP, using separate encoders for images and text trained with a contrastive loss. However, while CLIP (in our clip-vit-base-patch32 variant) uses a Vision Transformer for image encoding, ALIGN typically uses a CNN image encoder along with a Transformer for text. Moreover, the original ALIGN was trained on a much larger, noisier dataset (over 1.8 billion image–text pairs) than CLIP. In our experiments, we use an open-source version [^27] https://huggingface.co/kakaobrain/align-base that preserves the core architecture but is trained on different data, achieving even better accuracy on some benchmarks.
3. FLAVA [60]: [^28] https://huggingface.co/facebook/flava-full a unified multimodal model developed by Facebook that jointly learns representations for images and text. Unlike CLIP, which primarily relies on a contrastive approach applied to image-text pairs, and ALIGN, which emphasizes scaling up representation learning using noisy text supervision with separate encoders, FLAVA adopts a holistic pre-training strategy. It integrates both unimodal and multimodal objectives, enabling FLAVA to capture richer semantic interactions across modalities.
4. BLIP + MiniLM: This pipeline approach first converts images into descriptive captions using BLIP [35] [^29] https://huggingface.co/Salesforce/blip-image-captioning-base—a state-of-the-art image captioning model from Salesforce known for its high-quality, informative captions. The generated captions are then transformed into embeddings with MiniLM [72] [^30] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2, a robust text encoder widely adopted in industrial applications. This decoupled strategy not only leverages mature text retrieval systems but also enables independent optimization of the captioning and text-embedding stages. As a result, it has become a strong baseline and is employed in many commercial products, offering enhanced interpretability and scalability compared to end-to-end models like CLIP.
E.2 Dataset Details
((a)) Blue napkin to the right of white plate
((b)) White napkin to the left of blue plate
Figure 6: Illustration of COLA compositional benchmark queries, using each image description, where the goal is to identify the correct image
Object Detection Datasets:
  1. 1.
    Caltech256 [21]: Contains 30,607 images spanning 256 object categories. Its diverse object classes make it a useful benchmark for assessing the robustness of detection and retrieval models.
  2. 2.
    MS COCO [38]: With over 118K images and 80 object categories, MS COCO is widely used for object detection and segmentation. Its comprehensive annotations and varied scene compositions provide a challenging testbed.
  3. 3.
    LVIS [22]: offers instance segmentation with a long-tail distribution of over 1,200 categories. Its focus on rare and fine-grained objects is critical for evaluating retrieval performance on less frequent classes.
  4. 4.
    BDD100k [79]: This dataset is tailored to urban driving scenarios with detailed annotations, making it valuable for testing retrieval in real-world, dynamic contexts.
Natural Language Query Datasets:
For evaluating the performance of various baselines on complex natural language queries, we focus on scenarios where models must differentiate between two highly similar images that exhibit nuanced differences. In these settings, the model is expected to identify the correct image by carefully attending to subtle details in both the visual content and the corresponding query. To rigorously assess these capabilities, we utilize several compositional benchmarks:
  1. 5.
    COLA [56]: As a benchmark for compositional text-to-image retrieval, this dataset features nuanced captions that require the system to distinguish between multiple similar images by capturing subtle semantic and spatial details.
  2. 6.
    Winoground [63]: Designed to evaluate visio-linguistic alignment, Winoground presents pairs of images and captions that differ only in their compositional structure. This dataset challenges models to accurately map nuanced language to the corresponding image, serving as a stringent test of fine-grained retrieval performance.
  3. 7.
    NoCaps [2]: A large-scale benchmark for novel object captioning, featuring images from Open Images with human-generated captions. It is crucial for evaluating a model’s ability to handle zero-shot retrieval on images containing objects not commonly found in standard captioning datasets.
  4. 8.
    SentiCap [44]: Derived from datasets like MS COCO, SentiCap provides images with captions specifically annotated for positive or negative sentiment. It allows for evaluating retrieval systems’ ability to understand and retrieve images based on subjective or emotional language cues, moving beyond purely factual descriptions.
E.3 Evaluation Metrics
We use different evaluation metrics for object detection and complex natural language queries benchmarks. In object detection benchmarks, we define each benchmark dataset $\mathcal{D}$ as a set of images $\mathcal{D}=\{t_{1},t_{2},\ldots,t_{p}\}$, where each image $t_{i}$ comprises a set of objects $t_{i}=\{o_{1},o_{2},\ldots,o_{s}\}$, with $o_{j}$ representing an individual object within image $t_{i}$. For each benchmark, we compute the union of all objects present across all images, denoted as $\bigcup_{i=1}^{n}t_{i}=\bigcup_{i=1}^{n}\{o_{1},o_{2},\ldots,o_{s}\}$, and utilize each unique object $o_{j}$ within this union as a query $q$ for each retrieval engine. For a given query $q$, we retrieve a ranked list of files =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{n}\}$ ($n=60$), representing the retrieved images. We construct a list of relevance scores =$\mathcal{R}=\{r_{1},r_{2},\ldots,r_{n}\}$, where $r_{i}=1$ if $q$ is present in image $f_{i}$, and $r_{i}=0$ otherwise. We start list indexing from 1, so by definition $\mathcal{R}[j]=r_{j}$. We assume access to a function $C(.)$, which gets an object as input and returns the number of images that this object appears in. We are interested in retrieving at most 10 positive instances. Therefore, we define the effective number of positives as: $ep=\min(10,C(q))$.
Given $\mathcal{R}$ and $ep$, We assess the performance of the retrieval using the following metrics:
Mean Recall at $k$ (R@k): $r@k=\frac{\sum_{i=1}^{k}\mathcal{R}[i]}{ep}$. A higher r@k indicates that a larger proportion of the relevant images (up to 10) are retrieved in the top $k$ positions, we report mean r@k (R@k) which is the average of r@k over all objects in the dataset.
Precision at $k$ (P@k): $p@k=\frac{\sum_{i=1}^{k}\mathcal{R}[i]}{k}$. A higher p@k reflects that a greater proportion of the retrieved images are relevant, which is crucial when the user examines only the top results, same as before, we report the averaged out P@k over all object in the dataset.
Mean Average Precision (MAP): measures the average precision (AP) over all queries.
$AP=\frac{1}{ep}\sum_{i=1}^{n}\left(\frac{\sum_{j=1}^{i}\mathcal{R}[j]}{i}\right)\cdot r_{i},$
where $r_{i}=\mathcal{R}[i]$ and the effective number of positives is defined as $ep=\min(10,C(q))$. The overall mean average precision (MAP) is then the average of $AP$ over all objects in the dataset. MAP provides a single-figure measure that considers both the precision and recall across the entire ranked list, rewarding systems that rank relevant images higher.
Mean Reciprocal Rank (MRR): computes the reciprocal rank for each query and then averages these values over all queries. For a given query $q$, let $k_{q}$ be the smallest index such that $\mathcal{R}[k_{q}]=1$, indicating the rank of the first relevant image. The reciprocal rank for $q$ is defined as:
$RR(q)=\begin{cases}\frac{1}{k_{q}},&\text{if }\exists\,k\text{ with }\mathcal{R}[k]=1,\\[5.0pt] 0,&\text{otherwise.}\end{cases}$
The Mean Reciprocal Rank (MRR) is the average value over all queries. MRR emphasizes the importance of retrieving at least one relevant image as early as possible for ranking.
For the complex natural language query experiments, we evaluate our method using both MRR and Pairing Accuracy. In these experiments, the baseline is given two images, $t_{1}$ and $t_{2}$, along with two captions, $c_{1}$ and $c_{2}$, and each caption must be assigned to its corresponding image correctly. Pairing Accuracy is defined as the number of correct assignments divided by the total number of queries. Note that the pairing accuracy for a random baseline is 0.25% (chance of choosing the correct combination over all possible combinations).
E.4 Hyper-parameter Ablation Study
((a)) BDD